OpenStudy (anonymous):

How many solutions? 3x+2x=20 and y+9=8x

OpenStudy (anonymous):

@Hero?

hero (hero):

Is there supposed to be a y variable in the first equation?

OpenStudy (anonymous):

oh yeah sorry it is 3x + 2y = 20... Sorry

hero (hero):

3x + 2y = 20 y = 8x - 9 Use substitution to solve

OpenStudy (anonymous):

ok... uh I dont know x. What am I 'sposed to do?

OpenStudy (anonymous):

well no need to solve if you need number of solutions. Since they are the straight lines with different slopes it will intersect at only one point

hero (hero):

Substitute the expression for y into the first equation 3x + 2(8x - 9) = 20 Then solve for x, then solve for y.

hero (hero):

If you only need to know the number of solutions, then two lines always intersect at a single point.

OpenStudy (anonymous):

yesh ok thanks! Cant believe i missed that! :(